If α , β be the roots of x^2 + px + c = 0 , find the values of
i) α^2 + αβ + β^2
ii) α^3 + β^3
iii) α^4 + α^2β^2 + β^4

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
a^2 + ab + b^2 = (a + b)^2 – ab
 
 = p^2 – c
 
similarily find all otthers
 
Regards
Arun (askIITians forum expert)